Tidwell Industries has the following overhead costs and cost drivers. Direct labor hours are estimated at 100000 for the year.

    Activity Cost Pool         Cost Driver         Est. Overhead         Cost Driver Activity    
Ordering and Receiving Orders $   65000 500 orders
Machine Setup Setups 247500 450 setups
Machining Machine hours 1362500 125000 MH
Assembly Parts 1090000 1000000 parts
Inspection Inspections 245000 500 inspections


If overhead is applied using traditional costing based on direct labor hours, the overhead application rate is

a. $8.

b. $30.

c. $14.

d. $11.
Answer #1

Total overhead = 65000+247500+1362500+1090000+245000 = 3010000

Overhead rate = Estimated overhead/estimated labor hour = 3010000/100000 = 30.10

So answer is b) $30